Subject: 3.5-m Users Committee minutes 6/09/08
From: Bruce Gillespie
Submitted: Thu, 12 Jun 2008 14:34:46 -0600
Message number: 1106
(previous: 1105,
next: 1107
up: Index)
APO 3.5-m Users Committee Phonecon, 6/09/08
Attending: Suzanne Hawley, Bruce Gillespie, Scott Anderson, John
Bally, Jon Holtzman, Michael Strauss, Bill Ketzeback, and Remy
Indebetouw
Absent: Al Harper, Jon Fulbright, Mark Klaene
Minutes taken by Bruce Gillespie
**********************************
User feedback, comments from institutional representatives:
None of the committee members present had any substantive user issues
to report.
**********************************
Discussion of telescope/instruments report - Gillespie:
The report is given below. Bruce highlighted the site issues (unusual
weather, water-leak fix, etc.). John Bally updated the NIC-FPS status
by saying that the etalon is now at CU being prepared for cryo
testing. Also, based on discussions with UVa instrument engineers, a
second new fan-out board will be fabricated this summer in the hope of
further reducing detector read noise. Bill Ketzeback talked about the
ongoing development and testing of the telescope autofocus capability,
which the Observing Specialists are using in beta test mode. These on-
sky tests are yielding encouraging results, and more tuning and
refinements are in the works. Suzanne mentioned that the Q3 schedule
was posted yesterday, and she invites the users to check their
programs. Because of the summer shutdowns and high demand for
TripleSpec, the bright time in Q3 was especially oversubscribed, more
so than usual.
*************************
3.5-m Telescope, Instruments, and CIF Projects Highlights, 4/30/08
through 6/05/08
Mark Klaene & Bruce Gillespie
0) Overview
Operations continued to be largely routine through the period. Unusual
spring weather has caused several nights of lost observing due to
several episodes of high winds and elevated dust levels. The
surrounding forest remained closed by the USFS (since May 1st) due to
extreme fire danger, although this has not affected operations. Minor
amounts of rain and elevated humidity have somewhat reduced the fire
danger, and the wind and dust are starting to subside as we approach
the summer monsoons. However, we are also coming into the seasonal
pine-tree mating season, and telescope closure guidelines will be
strictly enforced because pollen is somewhat caustic to mirrors and
difficult to remove with CO2 snow cleaning procedures.
On site-related matters, we noticed that recent water consumption at
the site was still above normal despite earlier repair work, and
another underground water leak was subsequently located and fixed.
Site electric power was also relatively unstable, due to the high
winds that have caused fluctuations and outages in the county-wide
grid. We are also planning to repair the fuel leak in the diesel
backup generator later this month, although the generator is still
operational in the meantime.
1) Telescope
Tertiary rotation acted up again even after replacing its encoder. The
engineering staff is working on diagnosing the problem, however a
detailed inspection of the system will not occur until the summer
shutdown when we have the entire system apart. A telescope autofocus
script is being tested and refined; initial on-sky results look
promising.
2) Instruments
DIS, SPIcam, and the Echelle operated smoothly throughout the period.
The new fan-out board replacement for NIC-FPS and the etalon removal
work was completed. The dysfunctional etalon is now at CU for
evaluation. The new fan-out board did not eliminate all of the excess
detector read noise, only about a 40% reduction in noise was seen. A
CU team returned to the site to work on the noise problem some more.
It is now believed that additional modifications to the fan-out board
will be required, involving adding capacitors.
TripleSpec commissioning continues. Preliminary user documentation is
now available at the APO home page. The visiting instruments, GFP and
APOLLO, were both operational.
3) CIF projects
We made progress on the implementation of Agile field rotation and
migration to the TR2 mirror cell port. During initial bench testing
one of the Hall effect sensors was discovered to be defective. The
motor was subsequently sent back to the vendor for warranty repair.
The machined parts for moving Fastcam to its new port are complete and
ready for installation, which will free up the TR2 port for Agile.
The tertiary support/actuator upgrade is progressing. Design work is
still slightly behind schedule but remains on track for an August
installation. Ongoing cleanup of rotator axis controller software
continued.
4) Summer Shutdowns (Unchanged from last month)
There will be three shutdown periods this summer. Two 3-day shutdowns,
one in June and one in July, will be for secondary and axis drive
work. In August, we are planning a 17-day shutdown for the tertiary
support/actuator upgrade, as well as for M1 washing and several
additional maintenance tasks (e.g., annual instrument servicing).
**********************************
Narrow-band filters for Agile - Hawley:
Suzanne raised the issue of possible user interest in using narrow-
band filters with Agile. Normally, people will be using our stock 3 x
3 inch broad-band SDSS or Johnson filters with Agile, but there may be
Agile science applications that favor narrow-band filters (e.g.,
planetary occultations). Because of the optical layout of the
instrument, the stock 2 x 2 inch narrow-band filters at APO will cause
significant vignetting in Agile's focal plane, which will severely
limit the field of view. Suzanne suggested that if there are users
planning to use Agile with narrow-band filters, they should contact us
so that we can determine if we need to procure the larger narrow-band
filters.
ACTION: Users committee members and users should contact Suzanne and
Mark Klaene (slh at astro.washington.edu, mark at apo.nmsu.edu) if
they intend to use narrow-band filters for their future Agile science
programs.
**********************************
TripleSpec observing during Q3 - Hawley:
Suzanne reminded us that TripleSpec will be used in campaign mode
during Q3 on a shared-risk basis, and that the UVa instrument
developers will be on-call during the campaigns to help with
unexpected problems. TripleSpec user documentation and test data is
available through our website, as well as pipeline data reduction
software.
**********************************
SDSS-III update - Gillespie, Hawley:
Bruce reported on the recent NSF review of the SDSS-III proposal.
Although we will not likely hear the official results of the review
for a few weeks, apparently the review was "successful" from the NSF's
perspective. Once the NSF makes its funding decision, the funding
partners (NSF, Sloan, DoE, and MOU institutions) will likely try to
have a conference to assemble a joint funding plan for the survey.
**********************************
ACTION ITEMS:
[open from previous months]:
none
[new actions from this month]:
ACTION: Users committee members and users should contact Suzanne and
Mark Klaene (slh at astro.washington.edu, mark at apo.nmsu.edu) if
they intend to use narrow-band filters for their future Agile science
programs.
**********************************
New business - Hawley:
Although not on the agenda, Suzanne mentioned that she wanted the
committee to be aware that we have received a proposal for a new
visiting instrument. John Bally and Jian Ge are jointly planning to
build a very-high-resolution NIR spectrograph that uses a silicon
immersion grating. It is conceivable that the prototype instrument
could be placed at the 3.5-m telescope as early as this winter. As
well as being a complementary instrument for extra-solar planet RV
studies, it could support follow-up observations of APOGEE survey
targets. A white paper is being prepared for further discussion.
**********************************
Next meeting:
The next Users Committee phonecon will be on Monday, 4 Aug, at 8:30 AM
Pacific Time (we will skip the meeting in July). The agenda and other
materials will be sent to the committee members during the preceding
week.
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 1106 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO